4/d-7/x=5/a solve for X
I am not sure what to do, I can't figure it out!

Answer by faceoff57(108)

4/d-7/x=5/a
xad(4/d-7/x)=xad(5/a)
4xa-7ad=5xd
4xa-5xd=7ad
x(4a-5d)=7ad
x=7ad/4a-5d